# A Multi-Objective Evaluation Framework for Analyzing Utility-Fairness Trade-Offs in Machine Learning Systems

### v1

@@ III-A Use-cases
To formalize the proposed method, we evaluate three ML use-cases, which are based on two scenarios as black-box and white-box, typically found in the literature, exclusively from a deployment perspective. We explicitly assume that ML models are trained and one is only seeking to characterize their performance from a multi-objective perspective including the model’s utility and one to many fairness objectives.
The first type of scenario considers a “black-box” ML system $h_{1}(x)$ to provide binary outcomes for an input $x$ such that $\hat{y}=h_{1}(x)$ where $\hat{y}\in\{0,1\}$. To measure the approximate Pareto solution $S$, we assume the availability of a dataset $X_{test}$ that carries annotations for all considered objectives, i.e., the expected output of the classification $Y_{test}$, and demographic attributes $A_{test}$. Fig. 1 contains a representation of this scenario in two optimization dimensions as $\omega_{1}$ and $\omega_{2}$. As there is no tuning possibility to select the model ($\tau=\emptyset$), this test evaluates the solution in the deployed ML system as it is provided.
Figure 1: Scenario 1: System evaluation as a black-box test.
@@
The second scenario defines the evaluation in a “white-box” manner for an ML system $h_{2}(x)$ that is tunable over prediction scores (logits) as $\hat{y}=h_{2}(x)$ where $\hat{y}\in[0,1]$. Model selection in $S$ may be achieved via $\tau$ so that a set of non-dominated solutions filtered by this parameter is available for performance assessment of the given ML system by using $X_{test}$ alongside $Y_{test}$ and $A_{test}$. This scenario is illustrated in Fig. 2 for two optimization dimensions, $\omega_{1}$ and $\omega_{2}$.
@@
In the following, we demonstrate each combination of these two scenarios in three assessment based use cases. We work with two synthetically generated approximate PF solutions, $System1$ and $System2$, which represent different systems across the use cases, and evaluate them in a three dimensional setting to simulate an assessment in one direction as utility and two directions as fairness. Thus, we have an overall insight into the comparative trade-off performance of different ML systems by performing possible evaluation strategies commonly encountered during the assessment of ML systems.
Figure 2: Scenario 2: System evaluation as a white-box test.
III-A1 UC-1 - The comparative evaluation of two black-box systems
@@
In this use-case, $System1$ and $System2$ are considered in a black-box manner as illustrated in Fig. 1 to assess their comparative performance using the proposed evaluation framework. Since both systems are assumed to be black-box, we only have the model for each as provided and assess the trade-off performance without any tuning.
@@
III-A2 UC-2 - The comparative evaluation of one black-box system with one white-box system as a hybrid case
@@
This is the use-case where the comparative performance of $System1$ and $System2$ is evaluated in a hybrid manner by applying black-box and white-box scenarios. $System1$ is assumed to be deployed as it is without any tuning capability (black-box), and $System2$ can be modified to have different settings based on user preference (white-box).
III-A3 UC-3 - The comparative evaluation of two white-box systems
@@
This use-case considers the assessment of $System1$ and $System2$ as white-box by tuning them according to specified preferences. Thus, we demonstrate how the trade-off capacities of the systems can be assessed when tuning is feasible and how model selection is achieved by fully leveraging their capabilities.
@@
We perform simulations for these use-cases in Section III-D to exemplify them quantitatively so that it is clarified how the proposed evaluation framework can be applied for such different assessment strategies given ML systems in comparison. These simulations are based on the synthetically generated systems, $System1$ and $System2$, and exhibit the PF trend with non-dominated and dominated points as expected from the utility-fairness trade-off systems.
III-B MOO Based Performance Indicators
@@
Central to MOO is the concept of the Pareto Front (PF), which delineates the set of all Pareto optimal solutions. A solution is deemed Pareto optimal if no other solution can enhance one objective without degrading another. In this regard, solutions residing on the PF are referred to as non-dominated solutions. More formally, given two points $\mathbf{x},\mathbf{x^{\prime}}$ in the multidimensional solution space $\Omega$, the sample $\mathbf{x}$ is said to dominate $\mathbf{x^{\prime}}$ as $\mathbf{x}\prec\mathbf{x^{\prime}}$, if $\mathbf{x}$ is farther from the reference point $r\in R$ than $\mathbf{x^{\prime}}$ in any considered dimension of $\Omega$, and is strictly farther at least for one of them. We note $r$, a.k.a. nadir point, represents the worst possible outcome for an experiment. In a minimization problem, for example, $\mathbf{x}$ provides a smaller combined value for the target objective when compared to $\mathbf{x^{\prime}}$, see illustration in Fig. 3.
@@
Figure 3: Dominance in bi-objective 3a minimization and 3b maximization problems: $x^{\prime}$ is dominated by $x$ with respect to the reference point $r$.
@@
$\mathcal{P}:=\{\mathbf{x}\in\Omega|\neg\exists\mathbf{x^{\prime}}\in\Omega\quad\mbox{such that}\quad\mathbf{x^{\prime}}\succ\mathbf{x}\}$ (1)
Pareto optimal solutions are called the Pareto set and the image of the Pareto set constructs the Pareto Front (PF) [37]. We note that, in real-world problems, the PF is rarely achievable. We refer to suboptimal solutions approximating the PF as $S$ [31] as shown in Fig. 4. We propose expanding on this approach for evaluating ML systems under multiple fairness constraints. This approach is analogous to the analysis of receiver operating characteristics (ROC) or detection-error trade-off curves in classical ML.
Whereas interpretation of a solution set $S$ considering two optimization dimensions $(\omega_{1},\omega_{2})$ is straightforward [16], concurrent analysis of multiple fairness constraints is typically done [14, 38] as a single degree of fairness by treating equity performances in isolation from one another. In this type of analysis, the dependency/correlation between different fairness criteria is not considered and the evaluation remains oversimplified. However, in a multi-task setting, every objective may conflict with each other, as one may not be improved without deteriorating others. This dependency between objectives, as is also the case for multiple fairness criteria alongside utility, should be projected into one shared space so that the multiple degrees of evaluation may be achieved in a fused way. To address this, we propose to characterize the solution set $S$, representing an ML system using multiple criteria from MOO. These indicators will be assembled in an easy to interpret table and a plot. Qualitative analysis can still be carried out when the number of concurrently analyzed objectives is small ($N=2$) or when visual clutter is minimal in systems with $N>2$. In all cases, the proposed evaluation framework via a PF characteristic remains usable.
III-B1 The Performance Indicators
In the design of metrics for MOO, four complementary performance criteria are typically considered to analyze the PF optimality: convergence, diversity, convergence-diversity, and capacity (or cardinality) [37, 39]. Measuring strict convergence, which denotes the proximity of the solution to the true PF, is not often attainable, we therefore focus on the other 3 properties, and describe them next.
@@ III-B1 The Performance Indicators
The UD indicator [32] evaluates the deviation characteristic of the non-dominated solution distribution and is formulated as:
$UD(S,\sigma)=\frac{1}{1+D_{nc}(S,\sigma)}$ (2)
@@
$D_{nc}(S,\sigma)=\sqrt{\frac{1}{|X_{n}|-1}\sum_{i=1}^{|X_{n}|}\left(nc(x^{i},\sigma)-\mu_{nc(x,\sigma)}\right)^{2}}$ (3)
@@
$nc(x^{i},\sigma)=|\{x\in X_{n},\|x-x^{i}\|<\sigma\}|-1$ (4)
@@
$\sigma$ is the niche radius that is problem dependent and can be adjusted based on the distribution of the candidate solution in the space. $\mu_{nc(x,\sigma)}$ is the mean of the niche counts, $nc$, calculated as $\mu_{nc(x,\sigma)}=\frac{1}{|X_{n}|}\sum_{j=1}^{|X_{n}|}nc(x^{j},\sigma)$. The UD indicator is expected to be lower for a more uniform solution set. This indicator evaluates how uniform the solution set is spanned in the metric space based on an upper-bound distance, $\sigma$. For instance, a system with the lowest UD value among others exhibits the best performance as its solutions are the most uniformly distributed. Having a trade-off system with a lower UD value corresponds to a more uniformly spanned set of non-dominated points. This increases the likelihood of achieving a desired combination of utility with fairness in tuning, compared to a system with a higher UD. In this study, we apply a transformation to UD by taking its inverse as $\overline{UD}=1-UD$ to make it consistent with other indicators so that a higher value means more uniformity. Although the UD measures the coverage of the solution space by the candidate set, it fails to characterize PF as any type of uniformly distributed solution (whether Pareto optimal or not) may yield high performance in terms of this indicator.
@@
The OS indicator [30] assesses the spread of the solutions obtained by the trade-off system. For a minimization problem evaluated in $N$ different dimensions, this indicator is formulated as:
$OS(S,\mathcal{P})=\prod_{i=1}^{N}\left|\frac{\max\limits_{s\in S}s_{i}-\min\limits_{s\in S}s_{i}}{\max\limits_{p\in\mathcal{P}}p_{i}-\min\limits_{p\in\mathcal{P}}p_{i}}\right|$ (5)
@@
where the nominator and denominator are the absolute difference between the worst and best points for the candidate solution $S$ and Pareto optimal set $\mathcal{P}$, respectively. A higher OS value indicates a more widely spread solution. This indicator assesses how well the points from the candidate set spread towards the ideal of the optimal PF. For instance, a system with a higher OS score compared to others has more points close to the ideal point and fewer ones near the nadir (here, we can access the nadir and ideal points without having the exact PF solution, so there is no requirement to know the PF a priori). Having a higher OS value exhibits a more spread characteristic for non-dominated solutions, leading to an improvement in tuning performance for the trade-off system when the selection of models around the ideal point is expected. In this study, OS is in the range of $[0,1]$ and there is no transformation applied as it is scaled compatible with other indicators. Similarly to capacity and distribution, this measurement also fails to analyze Pareto optimality in a comprehensive manner as it only assesses the extreme cases without considering the entire PF space. In Fig. 5, both UD and OS indicators are exemplified in synthetically generated data. $System1$ is said to have more uniformity and spread over $System2$ as its points are more equally distributed, $UD_{System1}=0.54<UD_{System2}=0.64$, (Fig. 5a) and it has more non-dominated points near the ideal, $OS_{System1}=0.45>OS_{System2}=0.05$, (Fig. 5b).
@@
Figure 5: Diversity: (a) $System1$ (blue) provides solutions that are more uniformly distributed than $System2$ (red) and therefore has lower $UD$. (b) $System1$ (blue) better covers the extremes of the PF approximations and therefore has better spread (larger $OS$) than $System2$ (red).
@@
This measurement evaluates both convergence and diversity together so that the proximity (convergence) alongside the distribution/spread (diversity) of the candidate solution set are projected into a single scalar score. This unary metric measures the volume of the objective space covered by an approximation set, relying on a reference point for calculation. The HV takes distribution, spreading, and convergence into account at the same time, making it unique in this regard. Recognized for its distinctive properties, HV is Pareto-compliant, ensuring that any approximation set achieving maximum quality for a MOO contains all Pareto optimal solutions. The reference point can be simply attained by constructing a solution of worst objective function values. Given a minimization based MOO problem with two objectives, as shown in Fig. 6, it is expected to have solution sets with points that are in the best achievable state in the objective space. This should be the case even if the objectives are conflicting with each other. PF is one possible setting for such cases with non-dominated solutions. In Fig. 6, $x_{1}$ and $x_{2}$ are two non-dominated solutions drawn from the PF-like solution set (represented as a dashed curve) with one dominated solution, $x^{\prime}_{3}$. The performance of such a solution set may be evaluated by the Hypervolume (HV) indicator to analyze how optimal the set is in terms of convergence and diversity. By discarding the dominated solution $x^{\prime}_{3}$, which should not be part of an optimal solution set, the HV indicator is calculated as the union of two volumes constructed between each of the non-dominated solutions, $x_{1}$ and $x_{2}$, and the reference point $r$ that is chosen as one of the poorly performing solutions in the space. The HV formulation is then as follows:
$HV=vol_{1}\cup vol_{2}=VOL\left(\prod_{i=1}^{2}[x_{1}^{i},r^{i}]\cup\prod_{i=1}^{2}[x_{2}^{i},r^{i}]\right)$ (6)
The formulation in (6) may be generalized as [40]:
$HV(S)=VOL\left(\bigcup_{\begin{subarray}{c}x\in S\\ x\prec r\end{subarray}}\prod_{i=1}^{N}[x^{i},r^{i}]\right)$ (7)
In this study, HV is on the scale of $[0,1]$ as every solution in the metric space is represented by measurements between 00 and $1$. An illustrative example in Fig. 7 shows how two systems are evaluated in terms of HV. $System1$ occupies a larger volume in 2D space compared to $System2$ as it is further away from the reference point. HV reflects this situation with $HV_{System1}=0.55>HV_{System2}=0.21$.
@@
This measurement quantifies the number of non-dominated points in the candidate solution set. The Overall Nondominated Vector Generation (ONVG) and Overall Nondominated Vector Generation Ratio (ONVGR) are some known capacity indicators. ONVG, proposed in [33], is the number of non-dominated solutions, $X_{n}$, in the candidate solution set, $S$, and formulated as:
$ONVG(S)=|X_{n}|$ (8)
As similarly proposed in [33], ONVGR is the ratio of the non-dominated solution cardinality against the cardinality of $S$ and it’s defined as:
$ONVGR(S)=\left|\frac{X_{n}}{S}\right|$ (9)
Figure 7: Convergence-Diversity: $System1$ (blue) covers a larger volume of the metric space when compared to $System2$ (red), relative to the reference $r$ (top-right). $System1$ is therefore considered to be a better choice with respect to this criteria.
Both ONVG and ONVGR yield higher scores for solution sets with greater capacity. These capacity-based indicators do not provide an extensive analysis like convergence or diversity do; they are only used as auxiliary indicators when other measurements are not discriminative. For instance, we can select a system with a higher ONVG over other systems when the convergence and diversity are the same for all. Furthermore, they may help analyze the effectiveness of the optimization, as a higher number of non-dominated points compared to dominated ones is a good indicator of how well the objective is approximated. Having a larger number of non-dominated points may also improve tuning the trade-off system as the possibility of finding an expected combination of utility alongside fairness would increase due to more optimal solutions in the objective space. Similarly to UD, we also apply a transformation just for ONVG by normalizing over the maximum value of it for the systems as $\widehat{ONVG}=\frac{ONVG}{\max(ONVG)}$ to make it in the same range as others. On the other hand, ONVGR is in the range of $[0,1]$ with 00 and $1$ indicating the absence of the non-dominated and dominated solutions, respectively. However, these measurements fail to capture PF optimality as the number of solutions does not provide information about the Pareto characteristic. Fig. 8 highlights that $System1$ exhibits more capacity characteristic compared to $System2$ in terms of ONVG and ONVGR as it has more non-dominated solutions, $ONVG_{System1}=8>ONVG_{System2}=2$, and a bigger ratio on overall solutions, $ONVGR_{System1}=0.80>ONVGR_{System2}=0.66$.
III-C Radar Chart: Compact Visualization

### v2

@@ 3.1 Use Cases
To formalize the proposed method, we evaluate three ML use cases, which are based on two scenarios as black-box and white-box, typically found in the literature, exclusively from a deployment perspective. We explicitly assume that the ML models are trained and one is only seeking to characterize their performance from a multi-objective perspective including the model’s utility and one to many fairness objectives.
[…]
@@
The second scenario defines the evaluation in a “white-box” manner for an ML system $h_{2}(x)\in\mathcal{H}$ that is tunable over prediction scores (logits) as $\hat{y}=h_{2}(x)$ where $\hat{y}\in[0,1]$. Model selection in $S$ may be achieved via $\tau$ so that a set of non-dominated solutions filtered by this parameter is available for performance assessment of the given ML system by using $X_{test}$ alongside $Y_{test}$ and $A_{test}$. This scenario is illustrated in Fig. 2 for two optimization dimensions, $\omega_{1}$ and $\omega_{2}$.
@@
In the following, we demonstrate each combination of these two scenarios in three assessment based use cases. We work with two synthetically generated approximate PF solutions, System1 and System2, which represent different systems across the use cases, and evaluate them in a two-dimensional setting to simulate an assessment in one direction as utility and the other as fairness. Thus, we have an overall insight into the comparative trade-off performance of different ML systems by applying possible evaluation strategies commonly encountered during the assessment.
Figure 2: White-box system evaluation (Scenario 2).
3.1.1 UC-1 - The comparative evaluation of two black-box systems
@@
In this use case, System1 and System2 are considered in a black-box manner to assess their comparative performance using the proposed evaluation framework. Since both systems are assumed to be black-box, we only have the model for each as provided and assess the trade-off performance without any tuning.
@@
3.1.2 UC-2 - The comparative evaluation of one black-box system with one white-box system as a hybrid case
@@
This is the use case where the comparative performance of System1 and System2 is evaluated in a hybrid manner by applying black-box and white-box scenarios. System1 is assumed to be deployed as it is without any tuning capability (black-box), and System2 can be modified to have different settings based on user preference (white-box).
3.1.3 UC-3 - The comparative evaluation of two white-box systems
@@
This use case considers the assessment of System1 and System2 as white-box by tuning them according to specified preferences. Thus, we demonstrate how the trade-off capacities of the systems can be assessed when tuning is feasible and how model selection is achieved by fully leveraging their capabilities.
@@
We perform simulations for these use cases in Section 3.6 to exemplify them quantitatively so that it is clarified how the proposed evaluation framework can be applied for such different assessment strategies given ML systems in comparison. These simulations are based on the synthetically generated systems, System1 and System2, and exhibit the PF trend with non-dominated and dominated points as expected from the utility-fairness trade-off systems.
3.2 MOO Based Performance Indicators
@@
Central to MOO is the concept of the Pareto Front (PF), which delineates the set of all Pareto optimal solutions. A solution is deemed Pareto optimal if no other solution can enhance one objective without degrading another. In this regard, solutions residing on the PF are referred to as non-dominated solutions. More formally, given two points $\mathbf{x},\mathbf{x^{\prime}}$ in the multidimensional solution space $\Omega$ ($\mathbf{x},\mathbf{x^{\prime}}\in\Omega$), the sample $\mathbf{x}$ is said to dominate $\mathbf{x^{\prime}}$ ($\mathbf{x}\prec\mathbf{x^{\prime}}$), if $\mathbf{x}$ is no worse than $\mathbf{x^{\prime}}$ in all considered objective dimensions and is strictly better in at least one of them. Geometrically, this implies that $\mathbf{x}$ lies farther from the reference point $r\in R$, also known as the nadir point, which represents the worst possible outcome in the objective space. In a minimization problem, for example, $\mathbf{x}$ provides a smaller combined value for the target objective compared to $\mathbf{x^{\prime}}$, see the illustration in Fig. 3.
@@
Figure 3: Dominance in bi-objective minimization (a) and maximization (b) problems: $x^{\prime}$ is dominated by $x$ with respect to the reference point $r$.
@@
$\mathcal{P}:=\{\mathbf{x}\in\Omega\;|\;\nexists\mathbf{x^{\prime}}\in\Omega\quad\mbox{such that}\quad\mathbf{x^{\prime}}\prec\mathbf{x}\}$ (1)
Pareto optimal solutions are called the Pareto set and the image of the Pareto set constructs the Pareto Front (PF) (Audet et al., 2021). We note that, in real-world problems, the PF is rarely achievable. We refer to suboptimal solutions approximating the PF as $S$ (Zitzler et al., 2003) as shown in Fig. 4. We propose expanding on this approach for evaluating ML systems under multiple fairness constraints. This approach is analogous to the analysis of Receiver Operating Characteristic (ROC) or detection-error trade-off curves in classical ML.
Whereas interpretation of a solution set $S$ considering two optimization dimensions $(\omega_{1},\omega_{2})$ is straightforward (Little, 2023), concurrent analysis of multiple fairness constraints is typically done (Buolamwini and Gebru, 2018; Luo et al., 2024) as a single degree of fairness by treating equity performances in isolation from one another. In this type of analysis, the dependency/correlation between different fairness criteria is not considered and the evaluation remains oversimplified. However, in a multi-task setting, every objective may conflict with each other, as one may not be improved without deteriorating others. This dependency between objectives, as is also the case for multiple fairness criteria alongside utility, should be projected into one shared space so that the multiple degrees of evaluation may be achieved in a fused way. To address this, we propose to characterize the solution set $S$, representing an ML system using multiple criteria from MOO. These indicators will be assembled in an easy to interpret table and a plot. Qualitative analysis can still be carried out when the number of concurrently analyzed objectives is small ($N=2$) or when visual clutter is minimal in systems with $N>2$. In all cases, the proposed evaluation framework via a PF characteristic remains usable.
3.2.1 The Performance Indicators
In the design of metrics for MOO, four complementary performance criteria are typically considered to analyze the PF optimality: convergence, diversity, convergence-diversity, and capacity (or cardinality) (Audet et al., 2021; Jiang et al., 2014). Measuring strict convergence, which denotes the proximity of the solution to the true PF, is not often attainable; we therefore focus on the other three properties, and describe them next.
@@ 3.2.1 The Performance Indicators
The UD indicator (Tan et al., 2002) evaluates the deviation characteristic of the distribution for non-dominated solutions, denoted as $X_{n}$, and is formulated as:
$\mathit{UD}(S,\sigma)=\frac{1}{1+D_{\mathit{nc}}(S,\sigma)}$ (2)
@@
$D_{\mathit{nc}}(S,\sigma)=\sqrt{\frac{1}{|X_{n}|-1}\sum_{i=1}^{|X_{n}|}\left(\mathit{nc}(x^{i},\sigma)-\mu_{\mathit{nc}(x,\sigma)}\right)^{2}}$ (3)
@@
$\mathit{nc}(x^{i},\sigma)=|\{x\in X_{n}\mid\|x-x^{i}\|<\sigma\}|-1$ (4)
$\sigma$ is the niche radius that is problem dependent and can be adjusted based on the distribution of the candidate solution in the space. $\mu_{\mathit{nc}(x,\sigma)}$ is the mean of the niche counts, nc, and is defined as
$\mu_{\mathit{nc}(x,\sigma)}=\frac{1}{|X_{n}|}\sum_{j=1}^{|X_{n}|}\mathit{nc}(x^{j},\sigma)$ (5)
@@
The UD indicator is expected to be higher for a more uniform solution set. This indicator evaluates how uniform the solution set is spanned in the metric space based on an upper-bound distance, $\sigma$. For instance, a system with the highest UD value among others exhibits the best performance as its solutions are the most uniformly distributed. Having a trade-off system with a higher UD value corresponds to a more uniformly spanned set of non-dominated points. This increases the likelihood of achieving a desired combination of utility with fairness in tuning, compared to a system with a lower UD. Although the UD measures the coverage of the solution space by the candidate set, it fails to characterize PF as any type of uniformly distributed solution (whether Pareto optimal or not) may yield high performance in terms of this indicator.
@@
The OS indicator (Wu and Azarm, 2001) assesses the spread of the solutions obtained by the trade-off system. For a minimization problem evaluated in $N$ different dimensions, this indicator is formulated as:
$\mathit{OS}(S,\mathcal{P})=\prod_{i=1}^{N}\left|\frac{\max\limits_{s\in S}s_{i}-\min\limits_{s\in S}s_{i}}{\max\limits_{p\in\mathcal{P}}p_{i}-\min\limits_{p\in\mathcal{P}}p_{i}}\right|$ (6)
@@
where the nominator and denominator are the absolute difference between the worst and best points for the candidate solution $S$ and Pareto optimal set $\mathcal{P}$, respectively. A higher OS value indicates a more widely spread solution. This indicator assesses how well the points from the candidate set spread towards the ideal of the optimal PF. For instance, a system with a higher OS score compared to others has more points close to the ideal point and fewer ones near the nadir (here, we can access the nadir and ideal points without having the exact PF solution, so there is no requirement to know the PF a priori). Having a higher OS value exhibits a more spread characteristic for non-dominated solutions, leading to an improvement in tuning performance for the trade-off system when the selection of models around the ideal point is expected. In this study, OS is in the range of $[0,1]$ and there is no transformation applied as it is scaled compatible with other indicators. Similarly to distribution, this measurement also fails to analyze Pareto optimality in a comprehensive manner as it only assesses the extreme cases without considering the entire PF space. In Fig. 5, both UD and OS indicators are exemplified in synthetically generated data. System1 is said to have less uniformity but more spread than System2 as its points are more equally distributed, $\mathit{UD}_{\textit{System1}}=0.54<\mathit{UD}_{\textit{System2}}=0.64$, (Fig. 5a) and closer to the extreme points, $\mathit{OS}_{\textit{System1}}=0.45>\mathit{OS}_{\textit{System2}}=0.05$, (Fig. 5b).
@@
Figure 5: Diversity: (a) System1 (blue) provides solutions that are less uniformly distributed than System2 (red) and therefore has lower UD. (b) System1 (blue) better covers the extremes of the PF approximations and therefore has better spread (larger OS) than System2 (red).
During the study, we observed that OS can decrease drastically when any objective fails to cover the full extent of the true PF. This sensitivity arises from the multiplicative characteristic of OS as a small value in one dimension sharply reduces the final score. To smooth this behavior, we introduce the Average Spread (AS) as a less sensitive variant of OS. AS simply replaces the multiplicative operator with summation, and is defined as:
$\mathit{AS}(S,\mathcal{P})=\frac{1}{N}\sum_{i=1}^{N}\left|\frac{\max\limits_{s\in S}s_{i}-\min\limits_{s\in S}s_{i}}{\max\limits_{p\in\mathcal{P}}p_{i}-\min\limits_{p\in\mathcal{P}}p_{i}}\right|$ (7)
@@
This measurement evaluates both convergence and diversity together so that the proximity (convergence) alongside the distribution/spread (diversity) of the candidate solution set is projected into a single scalar score. This unary metric measures the volume of the objective space covered by an approximation set, relying on a reference point for calculation. The Hypervolume (HV) (Zitzler and Thiele, 1998) takes distribution, spreading, and convergence into account at the same time, making it unique in this regard. Recognized for its distinctive properties, HV is Pareto-compliant, ensuring that any approximation set achieving maximum quality for a MOO contains all Pareto optimal solutions. The reference point can be simply attained by constructing a solution of worst objective function values. Given a minimization based MOO problem with two objectives, as shown in Fig. 6, it is expected to have solution sets with points that are in the best achievable state in the objective space. This should be the case even if the objectives are conflicting with each other. PF is one possible setting for such cases with non-dominated solutions. In Fig. 6, $x_{1}$ and $x_{2}$ are two non-dominated solutions drawn from the PF-like solution set (represented as a dashed curve) with one dominated solution, $x^{\prime}_{3}$. The performance of such a solution set may be evaluated by the HV indicator to analyze how optimal the set is in terms of convergence and diversity. By discarding the dominated solution $x^{\prime}_{3}$, which should not be part of an optimal solution set, the HV indicator is calculated as the union of two volumes constructed between each of the non-dominated solutions, $x_{1}$ and $x_{2}$, and the reference point $r$ that is chosen as one of the poorly performing solutions in the space. The HV formulation is then as follows:
$\mathit{HV}=\mathit{vol}_{1}\cup\mathit{vol}_{2}=\mathit{VOL}\left(\prod_{i=1}^{2}[x_{1}^{i},r^{i}]\cup\prod_{i=1}^{2}[x_{2}^{i},r^{i}]\right)$ (8)
The formulation in (8) may be generalized as (Navon et al., 2020):
$\mathit{HV}(S)=\mathit{VOL}\left(\bigcup_{\begin{subarray}{c}x\in S\\ x\prec r\end{subarray}}\prod_{i=1}^{N}[x^{i},r^{i}]\right)$ (9)
In this study, HV is on the scale of $[0,1]$ as every point in the solution space is represented by measurements between 0 and $1$. An illustrative example in Fig. 7 shows how two systems are evaluated in terms of HV. System1 occupies a larger volume in 2D space compared to System2 as it is further away from the reference point. HV reflects this situation with $\mathit{HV}_{\textit{System1}}=0.55>\mathit{HV}_{\textit{System2}}=0.21$.
@@
This measurement quantifies the number of non-dominated points in the candidate solution set. The Overall Nondominated Vector Generation (ONVG) and Overall Nondominated Vector Generation Ratio (ONVGR) are commonly used capacity indicators. ONVG, proposed by Van Veldhuizen and Lamont (2000), is the number of non-dominated solutions, $X_{n}$, in the candidate solution set, $S$, and is formulated as:
$\mathit{ONVG}(S)=|X_{n}|$ (10)
As similarly proposed by Van Veldhuizen and Lamont (2000), ONVGR is the ratio of the non-dominated solution cardinality to that of $S$, and is defined as:
$\mathit{ONVGR}(S)=\left|\frac{X_{n}}{S}\right|$ (11)
Figure 7: Convergence-Diversity: System1 (blue) covers a larger volume of the solution space than System2 (red), relative to the reference $r$ (top-right). Thus, System1 performs better in terms of HV.
Both ONVG and ONVGR yield higher scores for solution sets with greater capacity. These capacity-based indicators do not provide an extensive analysis like convergence or diversity do; they are only used as auxiliary indicators when other measurements are not discriminative. For instance, we can select a system with a higher ONVG over other systems when the convergence and diversity are the same for all. Furthermore, they may help analyze the effectiveness of the optimization, as a higher number of non-dominated points compared to dominated ones is a good indicator of how well the objective is approximated. Having a larger number of non-dominated points may also improve tuning the trade-off system as the possibility of finding an expected combination of utility alongside fairness would increase due to more optimal solutions in the objective space. We apply a transformation for ONVG by normalizing over the maximum value of it for the systems as $\widehat{\mathit{ONVG}}=\frac{\mathit{ONVG}}{\max(\mathit{ONVG})}$ to make it in the same range as others. On the other hand, ONVGR is in the range of $[0,1]$ with 0 and $1$ indicating the absence of the non-dominated and dominated solutions, respectively. However, these measurements fail to capture PF optimality as the number of solutions does not provide information about the Pareto characteristic. Fig. 8 highlights that System1 exhibits a more capacity characteristic compared to System2 in terms of ONVG and ONVGR as it has more non-dominated solutions, $\mathit{ONVG}_{\textit{System1}}=8>\mathit{ONVG}_{\textit{System2}}=2$, and a bigger ratio on overall solutions, $\mathit{ONVGR}_{\textit{System1}}=0.80>\mathit{ONVGR}_{\textit{System2}}=0.66$.
3.3 Radar Chart: Compact Visualization